find the exact area of a triangle whose side lengths are the roots of
x^3-18x+105x-200

Answer by fractalier(6550)

If you factor this you will get
x^3-18x+105x-200 = (x-8)(x-5)(x-5)
Setting this equal to zero, yields side lengths of 5 and 8.
Since A = (1/2)bh = (1/2)(5)(8) = 20